You are bullish on RVH stock. The current market price is $50 per share, and you have $600,000 of your own to invest. You borrow an additional $400,000 from your broker at an interest rate of 3% per week and invest $1,000,000 in the stock. RVH pays no dividends.


a. Suppose the price of RVH stock falls immediately after your purchase. The maintenance margin is 30%. How low can the price of RVH stock fall before you receive a margin call?

b. Suppose a week has passed. What is your rate of return if the price of RVH stock has gone up by 20%? 


1
Expert's answer
2022-02-23T12:45:58-0500

(a). maintenance margin ="\frac{no of shared price-borrowed ammount}{no of shares \times{price}}"

30%="((\frac{100000}{50})price-{400000})" "\div(\frac{100000}{50\times{price}})"

"\frac{30}{100}" ="\frac{2000price-400000}{2000 price}"

=0.3"\times{2000price}" ="2000{price}"-400000

=(2000-600)"{price}" =400000

"{price}" ="\frac{400000}{14000}"

=$285.71

(b). you buy 20,000 shares for $1000000.

These shares increase by 20% over the next week, that is $200000. You pay interest of

0.03"\times" $400000=$12000. The rate of return will be ;

"\frac{200000-12000}{600000}" =0.31

=31%